7. B Co has recorded the following data in the two most recent period.

            Total Costs of Production            Volume of Production

                              K                                                 Units

                         13,500                                               700

                         18,300                                            1,100

What is the best estimate of the company’s fixed costs per period?
Answer #1

Answer = 5100

Workings

Variable Costs = Change in costs / Change in units

=(18300-13500) / (1100-700)

=4800/400 = 12

Variable costs at production quantity 700 units = 700 units * 12 = 8400

Fixed Costs =Total costs - Variable costs

= 13500 - 8400 = 5100

Checking

At production level of 1100 units,

variable costs = 1100 units * 12 = 13200

Fixed Costs = Total costs - Variable costs

= 18300 - 13200 =5100